### v1

# A Survey on Self-play Methods in Reinforcement Learning

Abstract
Self-play, characterized by agents’ interactions with copies or past versions of itself, has recently gained prominence in reinforcement learning. This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts. Then it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios. Finally, the survey highlights open challenges and future research directions in self-play. This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
Index Terms:
Self-play, reinforcement learning, game theory, multi-agent
I Introduction
Reinforcement learning (RL) represents a significant paradigm [1] within machine learning, concerned with the optimization of decision-making processes through interaction with an environment. It’s fundamentally modeled using a Markov decision process (MDP), a mathematical framework that describes an environment in terms of states, actions, transitions, and rewards. Within an MDP, agents operate by observing states, executing actions according to defined policies, receiving subsequent rewards, and transitioning to subsequent states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators [2]. This fusion of deep learning with RL has been instrumental in handling high-dimensional state spaces, contributing to breakthroughs in various complex tasks.
Moreover, the transition from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [3, 4, 5]. In MARL, the interdependence of agents’ actions introduces significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. By addressing issues such as non-stationarity and coordination, self-play offers an approach where an agent interacts with copies or past versions of itself [6, 7]. This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [8, 9, 10, 11], chess [10, 11], poker [12, 13], and video games [14, 15]. In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[8, 10].
Although some research takes a broad perspective through empirical game-theoretic analysis (EGTA) [16], it is important to note that there are relatively few comprehensive surveys focusing exclusively on self-play. Among these, some studies address the theoretical safety of self-play [17], while others develop an algorithmic framework for self-play that unfortunately does not accommodate the Policy-Space Response Oracle (PSRO) series of algorithms [18]. Furthermore, another study concentrates exclusively on PSRO [19]. Although these varied studies are valuable, they do not offer a perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to bridge this gap.
The survey is organized as follows. Sec. II introduces the background of self-play, including the RL framework and basic game theory concepts. Sec. III proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework, clarifying the self-play landscape. In Sec. IV, a comprehensive analysis is performed to illustrate how self-play is applied in various scenarios. Sec. V describes open problems in self-play and explores future research directions. Finally, Sec. VI concludes the survey on self-play.
II Preliminaries
In this section, we first introduce the framework of RL. Next, we present the basic game theory concepts and the typical evaluation metrics used in self-play.
II-A RL Framework
[…]
In the RL context, agents interact with the environment based on the subsequent protocol: At each discrete time step $t$, every agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ is the parameters. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment undergoes a transition from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to maximize the expected discount accumulated rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
II-B Game Theory Concepts
II-B1 (Im)Perfect Information and (In)Complete Information
In a game characterized by perfect information, only one player moves at a time. Each player has a comprehensive understanding of the current game state, the full history of moves that have been made, and all potential future developments. If these conditions are not met, the game is considered to have imperfect information [25, 26]. In a game of incomplete information, there exists at least one player who is unaware of the payoff of another player; otherwise, it is a game of complete information [27].
For instance, Go is a game of both perfect and complete information. Players have full awareness of the entire game structure, including all possible moves, and they can see every move made by their opponent as they take turns to act (perfect information). Furthermore, if the outcomes are considered binary, such as win or loss, the payoff for the players is known to both sides (complete information).
II-B2 Normal-Form and Extensive-Form
The normal-form and extensive-form are two different ways of representing games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ is pure strategy space of all players. A vector $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A pure strategy defines a specific and deterministic action for a player in a game, while a mixed strategy designates a probability distribution over the set of pure strategies, allowing randomized actions. A mixed strategy for the player $i$ is a probability distribution $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ is a probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$. If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game. If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game. If finite players (especially two players) are involved and each player has a finite set of strategies, a game in the normal-form can be directly depicted in a matrix.
Specifically, in two-player zero-sum symmetric normal-form games, the pure strategy space for both player 1 and player 2 is identical, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. As the utility function $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, for simplicity, we can use only one utility function $u$ such that $\forall\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. The evaluation matrix captures the outcomes of the game by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is chance and can be regarded as a special agent. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
The Prisoner’s Dilemma serves as a classic example to illustrate various concepts in game theory. In a modified version of the dilemma, the outcomes are as follows:
  • If one player confesses (C) and the other lies (L), the confessor will serve 1 year in jail, while the liar will serve 8 years.
  • If both players choose to confess, they will each serve 7 years behind bars.
  • If both players choose to lie, they will each serve only 2 years behind bars.
The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must decide simultaneously whether to confess or lie, without knowledge of the other’s choice. Games played in this manner are referred to as static games. The normal-form representation is particularly suitable for static games, as it captures the simultaneous nature of decision-making (as depicted in Fig. 1a). Another variant is the sequential Prisoner’s Dilemma, where the second player makes their decision with knowledge of the first player’s action. Games of this nature are called dynamic games. The extensive-form representation is well-suited for dynamic games, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig. 1d).
Furthermore, a normal-form representation can be transformed into an extensive-form representation. For instance, the transformation from Fig. 1a to Fig. 1b involves representing the simultaneous decisions in a tree structure. In this extensive representation, the dotted lines indicate that certain statuses belong to the same information set. Conversely, an extensive-form representation can also be transformed into a normal-form representation. For example, the transformation from Fig. 1d to Fig. 1c involves condensing the sequential decisions into a matrix format. In this case, the pure strategy space for player 2 is $2\times 2=4$, reflecting the need for player 2 to have a strategy for each of player 1’s possible actions (confess or lie). For instance, the notation for one of player 2’s pure strategies $C\rightarrow C,L\rightarrow C$ in Fig. 1c indicates that player 2 chooses to confess when he knows player 1 has chosen to confess and player 2 also chooses to confess when he knows player 1 has chosen to lie. Other symbols in Fig. 1c follow the same logic.
(a) Matrix Representation of simultaneous Prisoner’s Dilemma in normal-form.
(b) Game Tree Representation of simultaneous Prisoner’s Dilemma in extensive-form.
(c) Matrix Representation of sequential Prisoner’s Dilemma in normal-form.
(d) Game Tree Representation of sequential Prisoner’s Dilemma in extensive-form.
Figure 1: The example of Prisoner’s Dilemma.
Compared to normal-form games, extensive-form games introduce sequential decision-making, adding complexity to the game structure. Extensive-form games also have a close relationship with MGs. In MGs with simultaneous moves, agents’ actions are unknown to each other, creating various histories that are condensed into a single information set. The game’s utility is the sum of rewards discounted over time [28]. Beyond normal-form and extensive-form games, as well as MGs, the analysis of complex Markov or extensive-form games often employs a higher-level abstraction: the meta-game. The meta-game facilitates the exploration of policy learning within these games, focusing not on isolated actions but on broader strategies that arise from the game’s dynamics. In this advanced normal-form context, the policy population consists of the strategies currently employed by players. Meta-strategies are mixed strategies that assign probabilities over the policy population in the meta-game.
II-B3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes adhere to a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A quintessential example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others. In real-world settings, games exhibit complexities that extend beyond theoretical models. [29] argues that real-world games have two salient features: first, practice usually leads to performance improvements; and second, there are a plethora of qualitatively distinct strategies, each with unique advantages and disadvantages. In such games, the strategies form a geometric topology resembling a spinning top, where the vertical axis represents the performance of the strategy, and the radial axis represents the length of the longest cycle.
II-B4 Stage Game and Repeated Game
A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players. A famous example of a stage game is the Prisoner’s Dilemma. A repeated game is derived from a stage game that is played multiple times. Formally, a repeated game based on a stage game $\mathcal{G}$ is defined by playing $\mathcal{G}$ for $T$ periods, where $T$ can be finite or infinite. The strategies in a repeated game are history-contingent, meaning that they can depend on the entire sequence of past plays. It’s important to note that a stage game or a repeated game can be either represented in the normal-form or extensive-form.
II-B5 Nash Equilibrium
For simplicity, $\pi_{i}$ denotes the strategy of player $i$, and $\pi_{-i}$ denotes the strategies of all players other than player $i$. Given $\pi_{-i}$, player $i$’s best response (BR) is the strategy that maximizes player $i$’s payoff:
$BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$ (1)
A strategy $\pi_{i}^{*}$ is an $\bm{\epsilon}$-BR to strategies $\pi_{-i}$ if:
$u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$ (2)
where $\epsilon$ is a pre-specified threshold.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for every player $i$:
$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*}),\forall\pi_{i},$ (3)
meaning that no player can benefit by changing their strategy unilaterally, given the strategies of all other players. In other words, an NE is a situation where the strategy chosen by each player is a BR to the strategies chosen by all other players.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for every player $i$:
$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$ (4)
meaning that no player can increase their payoff by more than $\epsilon$ by unilaterally changing their strategy.
However, computing NE is generally intractable in complex games, leading some researchers to utilize $\alpha$-Rank [30] and Correlated Equilibrium (CE) [31] as alternatives. Additionally, some studies resort to Replicator Dynamics [32] as a method to analyze and understand the evolution of strategies within these games.
II-B6 Team Games
The framework of a two-player zero-sum game can be naturally extended to encompass team-based zero-sum games. Von Stengel and Koller analyzed zero-sum normal-form games involving a single team competing against an adversary [33]. In this type of team game, consider a team denoted by $\mathcal{T}=\{1,2,\cdots,n-1\}$. The player $n$ is the adversary $(\mathcal{D})$. In this kind of zero-sum normal-form team games, for any player $i,j\in\mathcal{T}$, the utility functions satisfy $u_{i}(\bm{\pi})=u_{j}(\bm{\pi})=u_{\mathcal{T}}(\bm{\pi})$ and $u_{\mathcal{D}}(\bm{\pi})=-(n-1)u_{\mathcal{T}}(\bm{\pi})$. A zero-sum single-team single-adversary normal-form game can also be extended to the domain of extensive games [34]. For any player $i,j\in\mathcal{T}$ and all terminal nodes $z\in Z$, the utility functions satisfy $u_{i}(z)=u_{j}(z)=u_{\mathcal{T}}(z)$ and $u_{\mathcal{D}}(z)=-(n-1)u_{\mathcal{T}}(z)$. Let $I_{\mathcal{T}}$ denote the information set defined as $\bigcup_{i\in\mathcal{T}}I_{i}$, and let $A_{\mathcal{T}}$ represent the set of actions accessible in the information sets within $I_{\mathcal{T}}$.
In scenarios where teammates are unable to coordinate their strategies, the team-maxmin equilibrium (TME) emerges as the most suitable solution concept [33]. We denote the collection of sequences of player $i$ by $Q_{i}$, representing the sequence-form actions undertaken by player $i$. The sequence-form strategy is encapsulated by a function $p_{i}:Q_{i}\rightarrow\mathbb{R}$, which maps each sequence $q\in Q_{i}$ to its associated probability of execution. Formally, the TME is articulated as:
$\arg\max_{p_{1},\ldots,p_{n-1}}\min_{p_{n}}u_{\mathcal{T}}\prod_{i=1}^{n}p_{i}.$ (5)
Similar to the arguments in normal-form games [33], it can also be inferred that in extensive-form games, a TME exists uniquely, barring any degeneracy and this TME aligns with the team’s utility maximization of the NE [34].
II-C Evaluation Metrics in Self-play
In this section, we introduce various self-play evaluation metrics, including NASHCONV (Section II-C1), Elo (Section II-C2), Glicko (Section II-C3), WHR (Section II-C4), and TrueSkill (Section II-C5). While NASHCONV measures the distance from Nash equilibrium, the other four metrics evaluate relative skill levels and are compared in Table I. It’s important to note that although numerous other evaluation metrics exist, the metrics highlighted here are among the most widely used in the field.
Table I: Comparison of Relative Skill Evaluation Metrics.
Elo Glicko WHR TrueSkill
Uncertainty Modeling $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Ratings At Any Time $\times$ $\times$ $\checkmark$ $\times$
Multiplayer In One Team $\times$ $\times$ $\times$ $\checkmark$
Bayesian Foundation $\times$ $\times$ $\checkmark$ $\checkmark$
II-C1 NASHCONV
Nash convergence (NASHCONV) serves as a metric to measure the deviation of a particular strategy from an NE. A lower NASHCONV value suggests that the strategy is closer to an NE, implying that no player would benefit from unilateral deviation from the strategy. Formally, it is defined as:
$\text{NASHCONV}(\bm{\pi})=\sum_{i}\max_{\pi_{i}\in\Pi_{i}}u_{i}(\pi_{i},\pi_{-i})-u_{i}(\bm{\pi}),$ (6)
where $\bm{\pi}$ denotes the combined strategy profile of all participating agents. In particular, in the context of two players, this deviation is commonly referred to as exploitability.
II-C2 Elo
The Elo system [35] operates under the assumption that the performance of each player in each game is a normally distributed random variable, with the mean being the player’s current rating. In a match between player A and player B, $R_{A}$ and $R_{B}$ are the current ratings of player A and player B. The probability density functions for the performance of player A and player B are given by:
$f_{A}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{A})^{2}}{2\sigma^{2}}},$ (7)
$f_{B}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{B})^{2}}{2\sigma^{2}}}.$ (8)
$E_{A}$ and $E_{B}$ denote the expected score (or the probability of winning) for player A and player B:
[…]
For convenience, the logistic function is used to approximate the probability (using the logistic curve with base 10 and choosing 400 to scale and normalize the difference in ratings):
$E_{A}$ $=\frac{1}{1+10^{(R_{B}-R_{A})/400}},$ (11)
$E_{B}$ $=\frac{1}{1+10^{(R_{A}-R_{B})/400}}.$ (12)
Note that $E_{A}+E_{B}=1$. $S_{A}$ is the actual outcome of the match for player A (1 for a win, 0.5 for a draw, 0 for a loss). $S_{B}=1-S_{A}$. After a match, the ratings are updated based on the difference between the actual outcome and the expected outcome. The adjustment is given by:
$\Delta R_{A}=K(S_{A}-E_{A}),$ (13)
$\Delta R_{B}=K(S_{B}-E_{B}).$ (14)
$K$ is a scaling factor, often determined by the specific domain of the application, and controls the maximum possible rating change for a single match. If two players have nearly identical ratings, the expected outcome will be close to 0.5. A win for either player will result in a moderate increase in their rating. In contrast, if there is a significant rating difference, the expected outcome will be heavily skewed. If the higher-rated player wins, their rating will increase by a value much smaller than $K$, reflecting the anticipated nature of their victory. However, if the lower-rated player secures an unexpected win, their rating will surge by a value approaching $K$, signifying the upset. The updated ratings become:
$R_{A}^{\prime}=R_{A}+\Delta R_{A},$ (15)
$R_{B}^{\prime}=R_{B}+\Delta R_{B}.$ (16)
However, there are challenges and limitations in Elo. First, the Elo system assumes that all matches are equally important. This might not be the case in all domains. Second, $K$ is often kept constant. However, a dynamic $K$ factor might be more appropriate, especially for players new to the rating pool or in scenarios where the importance of the match varies. Third, the standard Elo system does not incorporate a decay mechanism to account for the potential degradation or improvement of skills during periods of inactivity. Fourth, the basic Elo system is designed for one-on-one competitions. Adapting it to team or multi-player scenarios, such as team sports or online multiplayer games, can be challenging. Lastly, an important limitation of the Elo rating system is its unsuitability for games that exhibit high levels of non-transitivity [36].
II-C3 Glicko
The Glicko system refines the Elo system by introducing a measure of uncertainty or reliability in a player’s rating, termed rating deviation [37]. The primary motivation is to account for the variability in a player’s performance and the potential changes in skill over time. The Glicko-2 system, an extension of the original Glicko system, further refines these concepts and introduces the rating volatility $\sigma$, indicating the degree of expected fluctuation in the player’s rating [38].
In the Glicko-2 system, $r$ denotes the current rating of the player. $RD$ is the player’s rating deviation, and $\sigma$ is the player’s volatility. Convert the rating and rating deviation to the Glicko-2 scale:
$\mu$ $=\frac{r-1500}{173.7178},$ (17)
$\phi$ $=\frac{RD}{173.7178}.$ (18)
Then, calculate $v$ representing the estimated variance of the player’s rating based on game outcomes, $E(\mu,\mu_{j},\phi_{j})$ representing the probability of a player with rating $\mu$ defeating an opponent player $j$ and $\Delta$ representing the estimated improvement based only on game outcomes $s_{j}$:
$v$ $=\left[\sum_{j=1}^{m}g(\phi_{j})^{2}\{1-E(\mu,\mu_{j},\phi_{j})\}\right]^{-1},$ (19)
$E(\mu,\mu_{j},\phi_{j})$ $=\frac{1}{1+\exp(-g(\phi_{j})(\mu-\mu_{j}))},$ (20)
$\Delta$ $=v\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (21)
where:
$g(\phi)$ $=\frac{1}{\sqrt{1+3\phi^{2}/\pi^{2}}}.$ (22)
The update for $\sigma$ is more involved and requires an iterative procedure to solve for its new value $\sigma^{\prime}$. Then, calculate new $\mu^{\prime}$ and $\phi^{\prime}$:
$\mu^{\prime}$ $=\mu+\phi^{\prime 2}\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (23)
$\phi^{\prime}$ $=\frac{1}{\sqrt{\frac{1}{{\phi^{*}}^{2}}+\frac{1}{v}}},$ (24)
where:
$\phi^{*}=\sqrt{\phi^{2}+\sigma^{\prime 2}}.$ (25)
After calculations, the rating and rating deviation are converted back from the Glicko-2 scale:
$r^{\prime}$ $=173.7178\mu^{\prime}+1500,$ (26)
$RD^{\prime}$ $=173.7178\phi^{\prime}.$ (27)
II-C4 WHR
The Whole-History Rating (WHR) system [39] is a Bayesian rating system designed to estimate players’ skills from their entire game history. It is particularly adopted in handling the temporal dynamics of player skills. $R_{i}(t)$ denotes Elo rating of the player $i$ at time $t$. Similar to Equ. (11) and Equ. (12), $E_{A}(t)$ and $E_{B}(t)$ denote the expected score (or the probability of winning) for player A and player B at time $t$:
$E_{A}(t)$ $=\frac{1}{1+10^{(R_{B}(t)-R_{A}(t))/400}},$ (28)
$E_{B}(t)$ $=\frac{1}{1+10^{(R_{A}(t)-R_{B}(t))/400}}.$ (29)
Moreover, WHR assumes that ratings of each player $i$ is a Wiener process:
$r_{i}(t_{2})-r_{i}(t_{1})\sim\mathcal{N}(0,|t_{2}-t_{1}|\omega^{2}),$ (30)
where $\omega$ is a parameter representing the variability of ratings in time. Thus, if $r^{\prime}=R_{A}(t_{1}),r^{\prime\prime}=R_{A}(t_{2})$:
$p(r^{\prime\prime}|r^{\prime})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2}(\frac{r^{\prime\prime}-r^{\prime}}{\sigma})^{2}},$ (31)
where $\sigma=\omega\sqrt{|t_{2}-t_{1}|}$. In addition, a Wiener process is memoryless (or Markovian).
$\mathbf{R}(t)$ denotes player A’s ratings, $\mathbf{G}$ denotes the observation of game results. Thanks to the Bayes formula, the WHR system is to solve:
$\arg\max_{\mathbf{R}}p(\mathbf{R}|\mathbf{G})=\arg\max_{\mathbf{R}}\frac{p(\mathbf{G}|\mathbf{R})p(\mathbf{R})}{p(\mathbf{G})}.$ (32)
$P(\mathbf{G}|\mathbf{R})$ can be derived from the product of Equ. (28) and $p(\mathbf{R})$ can be derived from the product of Equ. (31). Utilizing the iterative process of Newton’s method, we can ascertain the solution to the given problem.
II-C5 TrueSkill
TrueSkill [40] is based on a probabilistic graphical model that employs Bayesian inference and adapts to multiple players in multiple teams. TrueSkill 2 [41] extends the original TrueSkill model with several enhancements by taking into account the experience of a player, the affiliation with a team, and some game-specific factors such as the number of kills. For simplicity, TrueSkill is introduced using the following scenario: team 1 has player 1, while team 2 has player 2 and player 3. Ultimately, team 1 defeated team 2. This is a simpler case than the one presented in the original paper [40]. It can be described in an undirected probabilistic graphic model (Fig. 2). The skill of each player, denoted by $s_{i}$, is represented by a Gaussian distribution $N(s_{i};\mu_{i},var_{i})$. The performance of each player, denoted by $p_{i}$, is also represented by a Gaussian distribution $N(p_{i};s_{i},\beta^{2})$. The performance of each team is denoted by $t_{i}$. The difference in team performance is denoted by $d$. If $d>\epsilon$, where $\epsilon$ is a small positive threshold, then team 1 beats team 2. The TrueSkill algorithm’s updating mechanism within the context of undirected probabilistic graphs uses the sum-product message-passing algorithm. This ensures that the skill estimates are refined iteratively, leading to accurate and reliable ratings for each player.
Figure 2: An undirected probabilistic graphic model example of TrueSkill.
III Algorithms
Based on existing self-play work [18, 42, 43, 44], we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. The framework is adept at handling multi-homogeneous-player general-sum games. It’s important to note that although homogeneous players represent a specific subset of heterogeneous players, the latter can be reformulated as the former by expanding the dimensions of the input vector, which essentially entails embedding agent identity information. Moreover, as players are homogeneous, we assume each player shares the same policy population.
To clarify our framework, we describe it in a more easily understandable way. All players share a common policy population with a fixed maximum size. In each iteration, a new policy is initialized for training, and opponent policies are sampled from the existing policy population. During this iteration, the opponent policies typically remain fixed, while only the policy being trained is updated. After the training process, the new policy replaces one of the policies in the policy population. An evaluation metric is then used to assess the performance of the updated policy population. Based on this performance, the strategy for sampling opponents is adjusted for the next iteration. This process is repeated. For a more detailed and precise description, please refer to Sec. III-A.
In addition, we divide self-play algorithms into four primary groups: traditional self-play algorithms (Sec. III-B), the PSRO series (Sec. III-C), the ongoing-training-based series (Sec. III-D), and the regret-minimization-based series (Sec. III-E). We analyze how these four categories align with our framework in their respective sections and introduce representative algorithms in each category. Moreover, in Sec. III-F, we discuss the differences among these four categories and their connections to our proposed framework. We also explain why our framework is more general than existing frameworks. Furthermore, we summarize the major elements of key algorithms within our framework for each category in Table II.
III-A Framework Definition
Algorithm 1 the Framework of Self-play
0:  $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$$\triangleright$ Policy population.
0:  $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$$\triangleright$ Interaction matrix.
0:  $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$$\triangleright$ MSS.
1:  for $e\in[[E]]$ do
2:     for $\sigma_{i}\in\Sigma$ do
3:        Initialize $\pi_{i}^{h}$
4:        $\pi_{i}^{h}\leftarrow\text{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
5:        $\mathcal{P}\leftarrow\text{EVAL}(\Pi)$triangleright Get policies’ performance.
6:        $\Sigma\leftarrow\mathcal{F}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
7:     end for
8:  end for
9:  return  $\Pi,\Sigma$
In Algo. 1, we define a unified framework of self-play based on [43, 42, 44, 18]. First, the input of the framework is defined as follows:
  • $\Pi$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which is determined by specific algorithms. We will further introduce this policy condition function in the following sections. For simplicity, we denote each policy $\pi_{i}(\cdot|h(i))$ by $\pi_{i}^{h}$. It’s important to note that $i$ refers to the $i$th policy in the policy population, not to a player. Additionally, the hyper-parameter $N$ denotes the policy population size of $\Pi$, not the number of game players. Furthermore, $\Pi$ can be initialized in two different ways. First, $\Pi$ can be considered as initialized with $N$ placeholder policies, meaning that there is no actual initialization of the policies in practice. Instead, each policy will be actually initialized in the training iterations (Line 3 in Algo. 1). We refer to this approach as the placeholder initialization. Second, $\Pi$ can be initialized with $N$ real policies, which may include random initialization or pretrained models. We refer to this approach as the actual initialization. A policy $\pi_{i}^{h}$ is considered an ineffective policy if it serves as a placeholder policy; otherwise, it is deemed an effective policy. Additionally, Table II provides a clearer illustration of the initialization of $\Pi$ and the expression of $h(i)$ across different categories of the main self-play algorithms.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix of the policy population. $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$. Namely, $\sigma_{i}$ illustrates how to sample opponent(s)’ policies against policy $i$. For instance, let $\sigma_{i}$ represent the probability of policies for each opponent. In this context, $C_{1}=N^{n-1}$, where n denotes the number of players. Alternatively, $\sigma_{i}$ can also be considered as sampling parameters within a sampling network. Especially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 3). It’s important to note that, unlike the original PSRO framework [42], $\sigma_{i}$ here is not the meta-strategy of policy $i$. Instead, in the two-player setting, if $\sigma_{i}$ in our framework is the opponent’s meta-strategy against policy i, our framework can reduce to the original PSRO framework.
  • $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$: A meta-strategy solver (MSS) $\mathcal{F}$ takes the performance matrix $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$ as its input and produces a new interaction matrix $\Sigma\in\mathbb{R}^{N\times C_{1}}$ as its output. $p_{i}$ is the performance of policy i. For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\Sigma$ is a square matrix ($C_{2}=N$). In addition, Table II concludes the MSS of representative self-play algorithms across these four categories.
Figure 3: When $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, we can consider various examples of conventional self-play algorithms: In the Top section, we define $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from [43] and this concept is initially proposed by [44].
Next, the core process within the framework is as follows:
  • For epoch $e\in[[E]]$ (Line 1 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the ineffective policy is likely to be chosen for training in each iteration and turned into an effective one, while effective policies remain unchanged. Conversely, if effective policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ accurately rechoflects the number of updates performed. Additionally, Table II summarizes the values of $E$ for different categories of self-play algorithms.
  • Initialize $\pi_{i}^{h}$ (Line 3 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. For instance, it may be initialized randomly, by leveraging pre-trained models [8, 15] or through a recently updated policy. We provide detailed descriptions of the initialization process for each algorithm series. Also, these are summarized in Table II.
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 4 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [45]. However, it often requires considerable computational effort. This computational demand can be a limitation, particularly in complex environments or with large action spaces. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, which can be calculated using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Additionally, other specially crafted ORACLEs are either tailored to new MSSes [46, 47] or introduced to enhance diversity [48, 49, 50].
    Some details of Algo. 23 and 4 need to be mentioned. $\bm{\pi_{opp}}$ denotes opponents’ policies. Moreover, in Algo. 2, off-policy RL algorithms [2, 51, 52, 53], typically require a replay buffer to gather samples; however, for simplicity, they are not explicitly included. Furthermore, Line 1 in Algo. 2 suggests that RL may not always yield an ABR that meets specific criteria. For instance, AlphaGo Zero [9] stipulates that the new policy must achieve a win rate exceeding 55 percent against its predecessor. Therefore, validation of the new policy is sometimes required. In addition, it is important to note that Algo. 3 is specifically devised for R-NaD [54], while Algo. 4 is expressly designed for a series based on regret minimization. They will be introduced in detail in Sec. III-C5 and Sec. III-E respectively.
  • $\text{EVAL}(\Pi)$ (Line 5 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to gauge the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$. Moreover, only effective policies will be evaluated to save on computation. Notably, if the MSS produces a constant matrix irrespective of the input, the evaluation step can be skipped to reduce computations.
Algorithm 2 Compute the Oracle in RL
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  while $\pi_{i}^{h}$ is not valid do
2:     for trajectory $\tau\in[[\tau_{max}]]$ do
3:        sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
4:        $\bm{\pi}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$
5:        $s_{0},\bm{o_{0}}\sim\rho$triangleright Initial state.
6:        for $t\in[[t_{max}]]$ do
7:           $\bm{a_{t}}\sim\bm{\pi}(\bm{o_{t}})$
8:           $s_{t+1},\bm{o_{t+1}}\sim P(s_{t},\bm{a_{t}})$
9:           $\bm{r_{t}}\leftarrow\bm{R}(s_{t},\bm{a_{t}})$
10:        end for
11:        $\pi_{i}^{h}\leftarrow\text{update}(\pi_{i}^{h})$$\triangleright$ Using RL algorithms
12:     end for
13:  end while
14:  return  $\pi_{i}^{h}$
Algorithm 3 Compute the Oracle in Evolution Theory
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
2:  $\bm{\pi_{\text{reg}}}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$$\triangleright$ Regularization policies.
3:  Transformed the reward according to $\bm{\pi_{\text{reg}}}$.
4:  $\pi_{i}^{h}\leftarrow$ Use replicator dynamics to play the reward-transformed game until convergence.
5:  return  $\pi_{i}^{h}$
Algorithm 4 Compute the Oracle in Regret Matching
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  for each player j do
2:     sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
3:     $\bm{\pi}=(\pi_{i}^{h}(j),\bm{\pi_{opp}^{h}}(-j))$
4:     Use regret matching to play the game and obtain new regret minimization information added to $h(i)$.
5:  end for
6:  return  $\pi_{i}^{h}$
Moving forward, we categorize self-play algorithms into four primary groups. We provide a detailed analysis of how each algorithm in these categories integrates and aligns with our proposed framework.
III-B Traditional Self-play Algorithms
Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input. These algorithms can start with agents training against their most recent version, helping to identify and exploit weaknesses. Additionally, other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies. In this section, we will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
III-B1 Integration into Our Framework
We can incorporate the traditional self-play algorithms into our proposed framework (Algo. 1), using the following settings. First, the policy population $\Pi$ utilizes placeholder initialization, meaning that initially, the policies are placeholders rather than actually initialized strategies. This initialization approach is used because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because in traditional self-play algorithms, only an ineffective policy is likely to be chosen for training in each iteration, thereby turning it into an effective policy. Here, the policy population size $N$ serves as the upper limit for the number of effective policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initiated randomly, which signifies starting the learning process from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, since the policies in traditional self-play algorithms are not conditioned, we simply set $h(i)=\emptyset$.
Next, we outline the traditional self-play schemes. For the sake of simplicity, we operate under the following assumption:
Assumption 1.
In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,\leavevmode\nobreak\ \forall i$.
Based on Assumption 1, we can further deduce the following important corollary:
Corollary 1.
In traditional self-play algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
Proof.
The policy population gradually increases over time. Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are ineffective policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
III-B2 Vanilla Self-play
In vanilla self-play [6], agents are trained by competing against their latest versions, ensuring consistent and aligned learning. The MSS of vanilla self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (33)
No matter what $\mathcal{P}$ is, the MSS produces the same interaction matrix, similar to the iterative refinement process of the policy. Although this MSS is simple, it is utilized in many further works, so we refer to it as vanilla MSS. Although vanilla self-play is effective in transitive games, it can lead the agent to cyclic learning patterns in non-transitive games like Rock-Paper-Scissors.
III-B3 Fictitious Self-play
Fictitious Play (FP) [55] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-play (FSP) [13] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-play (NFSP) [56] is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior. However, in more recent approaches [42], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (34)
In FSP, the MSS continues to generate a constant interaction matrix. Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
III-B4 $\delta$-uniform Self-play
$\delta$-uniform self-play is introduced by [7]. The hyper-parameter $\delta$ ranges between 00 and $1$ used to select the most recent $\delta$ (percentage) of policies for uniform sampling to generate opponent policies. When policy $i$ is in the training phase, following Corollary 1, only the preceding $i-1$ policies hold significance. As opponents for policy $i$, we select policies from the discrete uniform distribution $[\delta(i-1),i-1]$. When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized. Thus, we can get the MSS of $\delta$-uniform self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \delta(i-1)\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (35)
$f(i)=\max\{\lceil(1-\delta)(i-1)\rceil,1\},$ (36)
[…]
III-B5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-play (PFSP)  [15] leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$. The MSS of PFSP is given by Algo. 5.
Algorithm 5 the PFSP Meta-strategy Solver
1:  function $\mathcal{F}(\mathcal{P})$
2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i$
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
4:  return  $\Sigma$
The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can be assessed by other metrics as well. A similar MSS in  [14] can be utilized to allocate higher probabilities to strategies that perform better or are harder to defeat.
III-B6 Independent RL
Independent RL is a fully decentralized method in MARL. This simplifies the learning process and is useful in competitive or adversarial settings, while it may lead to suboptimal outcomes in situations requiring cooperation. Independent RL can be seen as a special case of self-play algorithms. In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$. The MSS of independent RL simplifies to an identity matrix:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ i=j\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (37)
The difference between independent RL and vanilla self-play lies in how the opponents’ policies are handled during training. In vanilla self-play, the opponents’ policies are often fixed, making the training process stationary. In contrast, in independent RL, the opponents’ policies evolve along with the policy being trained, resulting in a non-stationary training process. Moreover, if off-policy RL methods are used in independent RL, the samples collected can still be useful for training, even if they are generated by a different policy. This allows the agent to leverage past experiences more effectively and learn from a wider range of scenarios.
III-C PSRO Series of Algorithms
Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles. These oracles are policies that approximate optimal responses to the current meta-strategies of other agents. Additionally, PSRO employs EGTA to update meta-strategy distributions, thereby incorporating a level of exploration in policy selection to mitigate the risk of overfitting.
III-C1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize placeholder initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we simply set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any conditioning function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [42] in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria [45, 46, 47].
For simplicity, we also follow Assumption 1. Similar to traditional self-play algorithms, we can derive the Corollary 2 using a similar proof as Corollary 1.
Corollary 2.
In the PSRO series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
III-C2 Double Oracle
Double oracle (DO) [45] is traditionally only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 the NE-based Meta-strategy Solver
1:  function $\mathcal{F}(\mathcal{P})$
2:   $\sigma_{i+1}\leftarrow$ SOLVE-NASH$(\mathcal{P}_{1:i,1:i})$triangleright Opponent’s NE meta-strategy.
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
4:  return  $\Sigma$
III-C3 PSRO
PSRO [42] extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to a wide range of game scenarios.
III-C4 PSRO Variants
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. Through the establishment of a hierarchical structure of RL workers, Pipeline PSRO [57] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO [58] introduces the unrestricted-restricted (URR) game to narrow down the selection of opponent policies, thereby obviating the need for meta-game simulations in traditional PSRO. And, similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, XDO [59] allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, thereby enhancing its tractability in extensive-form games. ODO [60] integrates the no-regret analysis of online learning with the Double Oracle approach to improve both the convergence rate to an NE and the average payoff. Anytime PSRO [61] and Self-play PSRO [62] are designed to incorporate less exploitable policies into the policy population, thereby facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO [63] has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games [64, 65] and the selection problem in NE [66], Müller et al. put forward $\alpha$-PSRO [46]. Instead of NE, $\alpha$-rank [30], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, [47] proposes JPSRO to tackle multi-player general-sum games. It utilizes concepts of CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games. It also puts forward the CE and CCE-based MSS. Expanding on JPSRO, NeuPL-JPSRO [67] takes advantage of transfer learning. In addition to a shared parameter $\theta$, each strategy in NeuPL-JPSRO is characterized by a strategy embedding vector $v_{i}$. This approach avoids the need for training from scratch in each iteration, thereby enhancing efficiency.
Other research focuses on policy diversification, given that deriving a singular policy in highly transitive games often lacks significance. [48] introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. [49] uses determinantal point process [68] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO as well. Similarly,  [50] formulates behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO [69] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
III-C5 R-NaD
Although R-NaD [54] is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a special oracle computation technique (Algo. 3), which is executed in two stages: the first stage involves transforming the reward based on the regulation policy to make it policy-dependent, and the second stage applies replicator dynamics [32] until convergence to a fixed point. Crucially, in each iteration, the oracle added to the policy population $\Pi$ is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy will converge to the NE of the original game, provided that the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (33), the same as the MSS of vanilla self-play. This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
III-D Ongoing-training-based Series of Algorithms
In the PSRO series of algorithms, two key challenges arise. First, when operating with a limited budget, it is often necessary to truncate the ABR operators during each iteration. This can introduce sub-optimally trained responses into the population. Second, the redundant process of relearning basic skills in every iteration is not only inefficient, but it also becomes untenable when confronted with increasingly formidable opponents [43]. To address these challenges, the ongoing-training-based series of algorithms promote the ongoing training of all policies repeatedly. Namely, all effective policies are likely to be selected for training.
III-D1 Integration into Our Framework
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use actual initialization to initialize $\Pi$ because in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for a total of $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$. This means that policy updates are self-referential.
For the sake of simplicity, we also adopt Assumption 1. Different from Collory 1 and Collory 2, due to the continuous training process of all policies, we derive Collory 3.
Corollary 3.
In the ongoing-training-based series of algorithms, the interaction matrix $\Sigma$ is generally not a lower triangular matrix.
Proof.
When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an effective policy. Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning. As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
III-D2 FTW
Quake III Arena Capture the Flag is a renowned 3D multiplayer first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent [70] is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
$\mathcal{F}(\mathcal{P})_{ij}=\frac{1}{N}.$ (38)
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$. Furthermore, since these policies are not conditioned on any external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
III-D3 NeuPL
NeuPL [43] introduces another key innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$. Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
III-D4 Simplex-NeuPL
Simplex-NeuPL [71], which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess. To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex. Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(i))$ conditioned on the opponent sampling strategy $h(i)=\sigma_{i}$. Intriguingly, there is a slight possibility that $\sigma_{i}$ does not originate from the MSS. Instead, it is drawn as a sample from the population simplex. This sampling mechanism results in greater robustness.
III-E Regret-minimization-based Series of Algorithms
Another line of self-play algorithms is based on regret minimization. The key distinction between regret-minimization-based algorithms and the other categories is that they prioritize accumulated payoff over time, rather than focusing solely on a single episode. This approach leads to more aggressive and adaptable strategies, which are essential to avoid being exploited by opponents as time progresses. Additionally, these algorithms require players to deduce and adapt to opponents’ strategies over several rounds. This scenario is commonly observed in repeated games rather than stage games. For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game. It’s important to note that while traditional regret-minimization-based self-play doesn’t typically use RL, many subsequent research efforts have combined regret minimization with RL to achieve strong performance. In this section, we will also thoroughly discuss traditional regret-minimization-based methods to provide a foundation for understanding how integrating regret minimization with RL can lead to enhanced performance.
III-E1 Integration into Our Framework
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use placeholder initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy. Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy. More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$. Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. It’s important to note that the regret-minimization-based series relies heavily on the information contained in $h(i)$. For instance, in vanilla Counterfactual Regret Minimization (CFR) [72], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$ and once $h(i)$ is determined, the corresponding policy is also defined through regret matching. We will discuss vanilla CFR in detail in Sec. III-E2. Fifth, the ABR operator is described in Algo. 4, with the key point being to incorporate new regret minimization-based information into $h(i)$. Notably, while the original CFR updates regrets for all players simultaneously, this oracle (Algo. 4) updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. Namely, $h(i)$ is changing during the iteration $i$. This adjustment has not only been shown to empirically accelerate convergence but also possesses a theoretical error bound [73]. Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
Moreover, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (33). We can derive Collory 4.
Corollary 4.
In the regret-minimization-based series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix. More specifically, it is a unit lower shift matrix, with ones only on the subdiagonal and zeroes elsewhere.
Proof.
Regret minimization-based algorithms always use the latest strategy for training. In other words, in iteration $i$, $\pi_{i-1}^{h}$ is always chosen as the opponent policy. Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix. ∎
III-E2 Vanilla CFR
Regret measures the difference between the actual payoff and the best possible payoff that could have been achieved with a different strategy. The regret matching algorithm [74] in game theory optimizes decisions in iterative games by learning from past outcomes. More specifically, it involves selecting strategies based on accumulated positive overall regrets. Strategies with higher overall regret are generally more likely to be chosen, as players seek to correct past underperformance. After each round, players update the overall regret values for each strategy. When each player aims to reduce their average overall regret, their average strategies will converge towards an approximation of the NE over time. However, this traditional regret minimization algorithm applies primarily to normal-form games because computing overall regret in extensive-form games is challenging. Although extensive-form games can theoretically be converted into normal-form games, this conversion results in an exponential increase in states, rendering it impractical for complex extensive-form games.
[72] proposes counterfactual regret (CFR) minimization for extensive-form games with incomplete information. Here, it is referred to as vanilla CFR to distinguish it from subsequent advancements in the field. Vanilla CFR involves maintaining the strategy and the counterfactual regrets for each information set. Theoretically, the immediate counterfactual regret is specific to each information set, and the aggregation of these immediate counterfactual regrets can act as an upper bound for the overall regret. Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set. This simplification significantly reduces the computational complexity.
Next, there is a more detailed description of vanilla CFR. In this survey, $\pi_{i}$ is used to denote the strategy at iteration $i$, whereas the original paper uses $\sigma$. This change is implemented to maintain consistency throughout the discussion in this survey. Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$. The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
$v_{j}(\pi,I)=\sum_{h\in I,h^{\prime}\in Z}\text{Prob}_{-j}^{\pi}(h)\text{Prob}^{\pi}(h,h^{\prime})u_{j}(h^{\prime}).$ (39)
It is an unnormalized form of the counterfactual utility in the original paper. Based on this counterfactual value, the immediate counterfactual regret is defined as:
$R_{j,\text{imm}}^{T}(I)=\frac{1}{T}\max_{a\in A(I)}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (40)
where $\pi_{i}|_{I\to a}$ denotes player $j$ choose action $a$ with probability $1$ at iteration $i$. Moreover, the positive immediate counterfactual regret is:
$R_{j,\mathrm{imm}}^{T,+}(I)=\mathrm{max}(R_{j,\mathrm{imm}}^{T}(I),0).$ (41)
The player $j$’s average overall regret $R_{j}^{T}$ is defined as:
$R_{j}^{T}=\dfrac{1}{T}\max_{\pi(j)^{*}\in\Pi(j)}\sum_{i=1}^{T}\left(u_{j}(\pi(j)^{*},\pi_{i}(-j))-u_{j}(\pi_{i})\right).$ (42)
And theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
$R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (43)
Therefore, based on the theory discussed, the overall problem of regret minimization can be decomposed into numerous smaller regret minimization problems. This approach makes the problem manageable for extensive-form games of not excessively large sizes. In practical applications, the focus is primarily on immediate counterfactual regrets. For simplicity, we often drop "immediate" in discussions, thereby referring directly to counterfactual regrets. Consequently, the counterfactual regret associated with every action $a$ within every information set $I$ is recorded as follows:
$R_{j}^{T}(I,a)=\frac{1}{T}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (44)
$R_{j}^{T,+}(I,a)=\max\{R_{j}^{T}(I,a),0\}.$ (45)
The regret matching algorithm [74] is used to decide the strategy in each information set:
$\left.\pi_{j}^{T+1}(I)(a)=\left\{\begin{array}[]{ll}\frac{R_{j}^{T,+}(I,a)}{\sum_{a\in A(I)}R_{j}^{T,+}(I,a)}&\text{ if denominator}>0\\ \frac{1}{|A(I)|}&\text{ otherwise}\end{array}\right.\right..$ (46)
It’s worth noting that in some studies, the normalization factor $\frac{1}{T}$ is omitted in Equ. (40), (42) and (44).
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, which becomes computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$. These values are stored in $h(i)$ within our proposed framework (Algo. 1). This requirement leads to significant storage challenges.
III-E3 Time-saving Variants of CFR
Therefore, numerous studies focus on enhancing the time efficiency of CFR primarily through two main approaches. The first approach involves modifying the regret calculation to increase its speed. For example, CFR+ [75] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$. Note that $R_{j}^{+,i}(I,a)$ is computed slightly differently from $R_{j}^{i,+}(I,a)$ in Equ. (45). Additionally, CFR+ updates the regrets of each player sequentially. Furthermore, CFR+ adopts a weighted average strategy instead of a uniform average strategy, assigning a lower weight to the initial strategy for accelerated convergence. Beyond employing a weighted average strategy, [76] introduces the concept of weighted regrets and developed Discounted CFR (DCFR). Linear CFR (LCFR) is a specific case of DCFR, where LCFR assigns a linear weight of $i$ to regrets at iteration $i$.
The other approach involves adopting sampling methods. Although this approach requires more iterations, the duration of each iteration is reduced, ultimately decreasing the overall time to convergence. Monte Carlo CFR (MCCFR) [77] outlines a framework to incorporate sampling into CFR. This method divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for the calculation of sampled counterfactual values for each player, leading to immediate counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR represents a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, focusing on chance nodes alone. Moreover, [78] expands on chance-sampling into naive chance sampling, opponent/public chance sampling, self/public chance sampling, and public chance sampling. These methods differ in how they handle the sampling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [79, 80].
In addition to these two primary approaches, other studies have identified that warm starting [81] and pruning techniques [82, 83, 84] can accelerate CFR convergence as well.
III-E4 Space-saving Variants of CFR
Storing the strategies and regrets of each information set at each iteration requires substantial storage capacity. In games with perfect information, decomposition is employed to reduce the scale of problem-solving; that is, we solve subgames instead of the entire game. However, this approach faces challenges in imperfect information games. The difficulty arises in defining subgames, as they may intersect with the boundaries of information sets, complicating their delineation. CFR-D [85] is a pioneering method for decomposing imperfect information games with a theoretical guarantee. The game is divided into a main component, called the trunk, and several subgames. It posits that in an imperfect information game, a subgame can be conceptualized as a forest of trees that does not divide any information sets. In each iteration, CFR is applied within the trunk for both players, accompanied by the use of a solver to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root. Following these updates, the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving values $R^{i}(I_{*},a)$ for information sets located in the trunk and at each subgame’s root, which is denoted by $I_{*}$ at each iteration $i$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [12] and the Safe and Nested Subgame Solving technique used by Libratus [86]. We will discuss these approaches in Sec. IV-B1, where we explore their application to the specific context of Texas Hold’em.
III-E5 Estimation Variants of CFR
Although the above variants of CFR advance the field, they still fail to solve large imperfect-information extensive-form games directly. This limitation largely stems from the reliance on tabular representations for strategies. Typically, the process involves abstracting the original large game into a simpler form, applying the above CFR variants to this abstracted game, and then translating the developed strategies back to the original game. However, this abstraction process is highly game-specific and is heavily based on domain knowledge. Moreover, abstracting the game to a smaller scale often leads to suboptimal results compared to a more comprehensive abstraction, making the choice of abstraction scale critical for performance. Given these challenges, there is a shift towards estimation techniques. [87] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, the use of regression trees as the regressor limits RCFR’s applicability to smaller games, and the necessity for manually crafted features remains a drawback.
After advantage-based regret minimization (ARM) [88] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR [89] utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR [90] leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and its proven effectiveness in larger-scale poker games. Moreover, Single Deep CFR (SD-CFR) [91] demonstrates that training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM [92] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) [93] combines the thought of retrospective policy improvement with CFR.
III-F Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section and illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general. We also summarize the key points in Table II.
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and as training progresses, the policy population expands. Therefore, in our framework, we use placeholder initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes, designed to handle more intricate tasks. For example, $\alpha$-PSRO [46] specifically utilizes an $\alpha$-rank-based MSS to address multi-player general-sum games.
Unlike the two previously mentioned categories, the ongoing-training-based series employs a different paradigm by training the entire policy population together. This approach aims to strengthen all policies simultaneously at each epoch, rather than expanding the policy population and relying on newer policies to be stronger. As a result, actual initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential. Consequently, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
Lastly, the regret-minimization-based series focuses on the overall performance over time rather than on a single episode. For example, it is well-suited for games like Texas Hold’em, which require strategies involving deception, concealment, and bluffing. The main goal of the training process in this series is to update the regrets associated with different strategies. In our framework, we use $h(i)$ to store this information. Since the policies are determined by $h(i)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4). We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
Our framework is built upon PSRO [42] and NeuPL [43]. Here, we outline the differences between our framework and these two existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$, rather than being the meta-strategy of policy $i$. Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 4 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
Table II: Overview of Main Self-play Algorithms within Our Framework.
Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$
Vanilla Self-play [6] Equ. (33) $\emptyset$ Traditional Self-play $E=1$ Placeholder General
Fictitious Self-play [13] Equ. (34)
$\delta$-Uniform Self-play [7] Equ. (35)
PFSP [15] Algo. 5
Independent RL Equ. (37)
DO [45] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Placeholder General
PSRO [42] General
$\alpha$-PSRO [46] $\alpha$-rank-based
JPSRO [47] (C)CE-based
R-NaD [54] Equ. (33)
FTW [70] Equ. (38) $\emptyset$ Ongoing-training-based $E>1$ Actual $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$
NeuPL [43] General $\sigma_{i}$
Simplex-NeuPL [71] General $\sigma_{i}$
Vanilla CFR [72] Equ. (33) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Placeholder $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$
CFR+ [75] $R^{+,i}(I,a)$
CFR-D [85] $R^{i}(I_{*},a)$
RCFR[87] $\varphi(I,a)$
Deep CFR[87] $V(I,a|\theta_{p})$
IV Empirical Analysis
In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games and Mahjong, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play. We then illustrate how self-play is applied in each of these complex scenarios and provide a comparative analysis of these applications in Table III.
IV-A Board Games
The landscape of board games, the majority of which are perfect information games, was previously revolutionized by the introduction of two key techniques: position evaluation and Monte Carlo tree search (MCTS) [94, 95]. These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [96], checkers [97], othello [98], backgammon [99], and Scrabble [100]. In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level [101, 102, 103, 104, 105]. In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play. In addition to Go, we will broaden our exploration to include Stratego, a board game characterized by imperfect information, contrasting the majority of board games that are based on perfect information.
IV-A1 Go
The paradigm of Go is revolutionized with the launch of DeepMind’s AlphaGo series [8, 9, 10, 11], which leveraged the power of self-play to significantly elevate performance, setting a new benchmark in the field of Go.
In AlphaGo [8], the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS expansion and a precise policy network $p_{\sigma}(a|s)$. The second stage employs RL to refine the policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Self-play is instrumental here. More specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly chosen historical version $p_{\rho^{-}}(a|s)$, akin to the MSS shown in Equ. (34). Afterwards, $v_{\theta}(s)$ is trained using the samples from the games in which the trained policy network $p_{\rho}(a|s)$ competes against itself. In the final stage, MCTS integrates the policy and value networks to inform action selection.
Unlike AlphaGo, AlphaGo Zero [9] does not require any expert data except game rules and instead learns through self-play. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $p=\text{Pr}(a|s)$ and the state value $v$. Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, as opposed to the rollouts used in AlphaGo. Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (37). Furthermore, for a new policy to be incorporated into the policy pool, it must surpass a 55 percent win rate against its predecessor, aligning with the stipulations set in Algo. 2 at Line 1.
AlphaZero [10] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi, with some modifications. Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi. Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero.
Building upon AlphaZero, MuZero [11] takes the concept of learning from scratch to the next level, even operating without predefined game rules. MuZero incorporates ideas from model-based RL to model the dynamics of games. More concretely, in addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states. These three networks are trained jointly. Similarly to AlphaGo Zero and AlphaZero, MuZero employs MCTS guided by the three aforementioned networks to make decisions. The self-play process in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games. Furthermore, several studies have extended the capabilities of MuZero. For example, Stochastic MuZero [106] learns a stochastic model instead of a deterministic model to enhance its performance in more complex scenarios. Sampled MuZero [107] makes learning feasible in games with intricate action spaces.
IV-A2 Stratego
Unlike most board games which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and a large number of potential game states, estimated $10^{535}$ [108]. The game is divided into two phases: the deployment phase, where players secretly arrange their units, setting the stage for strategic depth, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [108] showed promising advances in AI’s ability to tackle it.
DeepNash scales up R-NaD [54] to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. In the dynamics stage, Neural Replicator Dynamics [109] is utilized to obtain the approximate fixed-point policy effectively. Additionally, DeepNash incorporates train-time fine-tuning and test-time post-processing methods to eliminate unreliable actions, thereby enhancing the robustness and accuracy of its output. DeepNash holds the third-place ranking among all Gravon Stratego players and secures victories in almost every match against existing Stratego bots.
IV-B Card Games and Mahjong
IV-B1 Texas Hold’em
Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements. In its two-player variant, it’s referred to as heads-up Texas Hold’em. The game becomes more complex with additional players. The gameplay begins with each player receiving two private cards (hole cards), followed by a round of betting. Subsequently, three community cards (the flop) are revealed, leading to another betting round. This is followed by the dealing of a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds. The objective is to construct the best five-card poker hand from any combination of hole cards and community cards. The betting proceeds until a showdown occurs, where the remaining players disclose their cards. The individual with the superior hand claims the pot. Moreover, Texas Hold’em is distinguished by its three betting formats: no-limit, fixed-limit, and pot-limit. Each format offers a different risk profile and complexity in terms of game state sizes. Among these, the no-limit format is particularly noted for its complexity, allowing players to bet any amount up to their entire stack of chips. Its popularity is partly due to its frequent appearance in tournaments and media, making it a staple in modern poker culture. While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis and educational contexts within game theory, our focus will focus on the development of algorithms designed to compete in real-world Texas Hold’em.
In poker, especially Texas Hold’em, abstraction is crucial to simplifying the inherent complexity of the game [110]. This methodology reduces the vast state space of the game to a manageable size, thus allowing algorithms to calculate strategies more efficiently. There are two main types of abstraction in Texas Hold’em: action abstraction, which simplifies the range of possible moves, and information abstraction, which groups similar hand strengths or pot sizes, streamlining the decision-making process.
Although heads-up limit Texas Hold’em (HULHE), the simplest form of Texas Hold’em with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [111]. It utilizes fixed-point arithmetic with compression and CFR+ [75] to address the issues of storage and computation, respectively, thus solving HULHE without involving any neural networks. This resulted in superhuman performance and surpassed previous agents. Moreover, NSFP [56] introduces the use of end-to-end RL through self-play training, achieving competitive performance in HULHE. Similarly, Poker-CNN [112], which utilizes a 3D-vector representation of HULHE combined with convolutional networks, learned through self-play and also achieved competitive performance. Deep CFR [90], combines deep neural networks with CFR to surpass the performance of NFSP. However, it’s noteworthy that despite advancements, network-based methods haven’t surpassed the regret-based method used by Cepheus, despite its substantial computational demands.
After the solution of HULHE, research focus shifted to heads-up no-limit Texas Hold’em (HUNL), representing a more complex challenge due to its significantly larger space of game states, approximately $10^{164}$ [113]. Thus, it is impossible to traverse the game tree as Cepheus does in HULHE. The primary technique employed by DeepStack [12] is continual re-solving. This method involves dynamically computing the strategy in real-time when it is the program’s turn to act, focusing on a subtree of limited depth and breadth. To estimate the outcomes of the furthest reaches of this subtree, DeepStack utilizes neural networks. These networks take into account both the player’s range and the opponent’s counterfactual values, which are monitored throughout the entire game. The neural networks employed include the turn network, the flop network, and the auxiliary network, all of which are pre-trained using 10 million randomly generated poker games. The target counterfactual values for these generated poker situations are derived through thousands of iterations using CFR+ [75]. DeepStack has demonstrated competitive performance against professional poker players by integrating these approaches. Libratus [86] employs self-play to develop its blueprint strategy based on an abstraction of the game, leveraging an enhanced version of MCCFR [77]. Initially, the blueprint strategy is deployed for decision-making in the first two rounds, benefiting from the detailed abstraction. In the later rounds, the blueprint strategy aids in nested safe subgame solving, akin to DeepStack’s continual re-solving aiming to get a more precise strategy for subgames in real-time. Additionally, Libratus features a self-improvement component designed to enhance the blueprint strategy by addressing any missing branches. Through these strategies, Libratus has proven capable of defeating professional HUNL players. ReBel [114] expands the concept of a game state in poker into a public belief state (PBS), thereby transforming the imperfect-information game into a perfect-information game with continuous state space. ReBel adopts the methodology of AlphaZero [10], which combines RL with a search algorithm to train both the value network and the policy network by self-play. To mitigate heavy computation and storage demands, AlphaHoldem [115] employs an end-to-end self-play RL method augmented with several additional techniques, to outperform its reimplementation of DeepStack. During training, the system employs the $K$-best Self-play, where it selects the top $K$ agents based on their ELO ratings [35] and the newly initialized agent. It then generates experience replays from competitions among these $K+1$ agents.
When there are more than two players in Texas Hold’em, the game becomes significantly more complex. Pluribus [116] based on the framework of Libratus [86], addresses this complexity in six-player no-limit Texas Hold’em. It employs both action abstraction and information abstraction to simplify the intricate dynamics of the original game. Similarly to Libratus, Pluribus utilizes MCCFR [77] in self-play to develop its blueprint strategy. Building on this blueprint strategy, it conducts a depth-limited search before executing actions. More specifically, diverging from Libratus, Pluribus maintains a streamlined policy pool consisting of only four strategies. This approach assumes that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently. Despite its strategy not adapting to the observed tendencies of its opponents, and lacking theoretical proofs of achieving an NE in six-player scenarios, Pluribus has demonstrated its capability by outperforming top professional players in six-player no-limit Texas Hold’em.
IV-B2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a strategic Chinese card game for three players. In this game, one player takes on the role of the landlord and competes against the other two players, who are the peasants. These two roles have distinct objectives: the landlord’s goal is to outplay the peasants, whereas the peasants work together to defeat the landlord. The essence of the game lies in its blend of strategy, cooperation among peasants, and competition between the two factions. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord through a process that involves assessing the strength of their cards. The outcome of this stage determines the game’s dynamics, with the landlord playing against the peasants. The card play stage is where the main action unfolds. Players take turns playing cards in various combinations, intending to be the first to empty their hands. The game’s complexity arises from its requirement for strategic thinking, both in playing cards effectively and in the peasants’ need to cooperate seamlessly. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played, adding a layer of unpredictability and strategic depth to the game. Moreover, with an estimated $10^{76}\sim 10^{106}$ possible game states [117] and an action space comprising $27472$ possible moves [118], DouDiZhu presents a vast landscape of strategic possibilities.
DeltaDou [119], built upon the AlphaZero framework [10], is the first algorithm to achieve expert-level performance in DouDizhu although it only considers the card play stage. It enhances the original framework by incorporating asymmetric MCTS to improve exploration in games characterized by imperfect information. Additionally, DeltaDou employs policy-value networks specially designed with input and output encodings to accelerate the training process. A notable feature of DeltaDou is its ability to infer the cards of other players using Bayesian methods with the help of the policy-value networks. DouZero [120] reduces training costs by opting for a sampling method over the tree search approach. It utilizes Monte Carlo methods combined with deep neural networks. More specifically, it works similarly to independent RL self-play to generate data, which is then stored in buffers. Subsequently, data is sampled from the buffers to update the neural networks. Remarkably, DouZero surpasses DeltaDou in performance, demonstrating the effectiveness of its methodology in mastering. Additionally, DouZero employs supervised learning to train a bidding network used during the bidding stage, utilizing expert data for this purpose. Based on DouZero, DouZero+ [121] enhances the original algorithm by incorporating opponent modeling, which assists the agent in making more informed decisions. Additionally, DouZero+ introduces coach-guided learning, aiming to ensure a balanced intensity of initial hand cards among the three players to further accelerate the training process. PerfectDou [118] leverages perfect information distillation within a Perfect-Training-Imperfect-Execution (PTIE) framework. In this approach, the value network is trained using both perfect and imperfect features, while the policy network is trained solely on imperfect features. The networks are refined through independent RL self-play, utilizing Proximal Policy Optimization (PPO) [122] with Generalized Advantage Estimation (GAE) [123]. Notably, PerfectDou does not incorporate the bidding stage into its model, thereby eliminating the need for expert data in this phase. PerfectDou achieves better performance than DouZero.
IV-B3 Mahjong
Mahjong, renowned for its complex interplay of skill, strategy, and chance, has evolved into various global variants, including the popular Japanese version known as Riichi Mahjong. This game is typically played by four participants who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges for artificial intelligence. Despite ongoing research, many AI implementations have yet to reach expert human levels [124, 125, 126]. AI systems must make decisions based on incomplete information and adapt strategies dynamically in response to the actions of multiple opponents, which complicates the decision-making process. Additionally, the game’s intricate winning rules and the enormous number of possible game states, estimated at around $10^{169}$ [117], add layers of complexity. These factors make Mahjong a particularly challenging domain for AI development, requiring sophisticated approaches to handle its unique demands effectively.
Suphx [127] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, specifically 10 dan on Tenhou [128], the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through self-play RL, employing several innovative techniques to enhance its effectiveness. These techniques include global reward prediction, which provides a continuous reward throughout the game, and oracle guiding, which offers perfect information at the start of self-play to guide strategy development. Additionally, parametric Monte Carlo policy adaptation is used to refine the policy further. Similarly, NAGA [129], developed by Dwango Media Village, and LuckyJ [130], designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players,
IV-C Video Games
In contrast to traditional board games and card games, video games often feature real-time actions, long time horizons and a higher level of complexity due to the broader range of possible actions and observations. Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
IV-C1 StarCraft II
StarCraft is a real-time strategy (RTS) game developed and published by Blizzard Entertainment. It showcases three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay experience. Renowned for its balanced gameplay, strategic depth, and challenging multiplayer features, the game tasks players with gathering resources, constructing bases, and building armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings.
AlphaStar [15], a significant advancement in AI, dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework bears similarities to that of AlphaGo [8], utilizing supervised learning initially to train the policy with data from human experts. Subsequently, it uses end-to-end RL and a hierarchical self-play method to further train the networks. More specifically, it divides all the agents into three types: main agents, league exploiters, and main exploiters. Additionally, it maintains a policy pool of past players that records all these types of agents. Main agents engage in both FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters use PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness and are added to the pool after achieving a high win rate or certain training steps, and are reset upon each addition. Most importantly, among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the training computation for AlphaStar is extensive. Further studies [131, 132, 133] have enhanced the league self-play training procedure by introducing innovative techniques to reduce computations.
IV-C2 MOBA Games
Multiplayer Online Battle Arena games (MOBA) are a popular genre of video games that blend RTS with role-playing elements. In typical MOBA games, two teams of players control their own unique characters, known as heroes and compete to destroy the opposing team’s main structure, often called the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map are key aspects of gameplay. Popular examples of MOBA games include League of Legends, Dota 2 and Honor of Kings. These games are known for their complex strategic depth, decision-making under conditions of imperfect information and emphasis on teamwork and player skill.
OpenAI Five [14] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. It employs distributed RL using PPO [122] along with GAE [123] to scale up training. Each player on a team shares the same policy network and receives shared observation inputs, with the only difference being a unique identity input for each player. The training process utilizes self-play, with an 80% chance of engaging in vanilla self-play and a 20% chance of using a self-play technique similar to PFSP used in AlphaStar [15]. Here, the probabilities of selecting each policy from the policy pool vary according to its quality score, which is continually updated based on competition results throughout the training. Higher quality scores increase the likelihood of a policy being selected. A new agent is added to the pool every 10 iterations. Similarly to AlphaStar, OpenAI Five requires extensive training resources. Thus, they developed surgery tools to resume training with minimal performance loss to manage continual changes in the environment and codes during development. Additionally, during training, OpenAI Five employs TrueSkill [40] to evaluate agent performance.
Another notable MOBA game, especially popular in China, is Honor of Kings. The 1v1 mode is conquered by [134], which boasts a significant winning rate against top professional players. This success can be attributed to well-designed systems tailored for large-scale RL training and sophisticated algorithm designs that handle complex control challenges effectively. A key element of its RL training methodology is the use of $\delta$-uniform Self-play. The 5v5 mode was later mastered by [135] as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, which exponentially increases the possible combinations of lineups and poses a learning collapse problem in RL training. It proposes curriculum self-play learning (CSPL) to mitigate this issue. Specifically, the training process is divided into three stages, drawing on insights from curriculum learning. The first stage involves training fixed lineups through self-play and utilizing human data to keep the two teams balanced to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for self-play with randomly picked lineups. This approach defeats professional player teams. Additionally, the self-play generated data is used to learn effective lineup drafting by utilizing MCTS and neural networks [136].
IV-C3 Google Research Football
Google Research Football (GRF) [137] is an open-source football simulator that emphasizes high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play, as opposed to testing on specially designed tasks. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. Additionally, it features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards, all of which contribute to its complexity.
WeKick [138] claimed victory in the GRF competition on Kaggle [139], which modified the game dynamics by allowing competitors to control just one player, either the ball carrier on offense or the nearest defender on defense. WeKick employed self-play strategies similar to those used in league training [15]. It initializes its opponent policy pool using strategies developed through reward shaping via RL, and through Generative Adversarial Imitation Learning [140], which facilitated learning from the tactics of other teams.
Further research delves into the full football game rather than controlling only one player in the team. Team-PSRO [141] extends PSRO [42] to team games, demonstrating approximate TMECor [34] and outperforming self-play in the 4v4 version of the full GRF. In the context of the 11v11 version of the full GRF, where the goalkeeper is rule-based controlled, TiKick [142] employs imitation learning to glean insights from data gathered by WeKick’s self-play, resulting in a multi-agent AI through distributed offline RL. Moreover, Fictitious Cross-Play (FXP) [143] proposes two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick in the 11v11 version of the full GRF. TiZero [144], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [15] to avoid expert data reliance, achieving a higher TrueSkill rating [40] than TiKick.
Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data
Players Information Game States
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
Go 2 $\checkmark$ $10^{360}$ AlphaGo [8] FSP $\checkmark$ $\checkmark$
AlphaGo Zero [9] Independent RL $\checkmark$ $\times$
AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
Stratego 2 $\times$ $10^{535}$ DeepNash [108] R-NaD $\times$ $\times$
Card Game HULHE 2 $\times$ $10^{17}$ Cepheus [111] CFR+ $\checkmark$ $\times$
NFSP [56] NFSP $\times$ $\times$
HUNL 2 $\times$ $10^{164}$ DeepStack [12] CFR+ $\checkmark$ $\times$
Libratus [86] MCCFR $\checkmark$ $\times$
ReBel [114] CFR-D / FSP $\checkmark$ $\times$
Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus [116] MCCFR $\checkmark$ $\times$
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [119] Independent RL $\checkmark$ $\times$
DouZero [120] Independent RL $\times$ $\checkmark$
PerfectDou [118] Independent RL $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx [127] Independent RL $\checkmark$ $\checkmark$
Video Game StarCraft II 2 $\times$ / AlphaStar [15] FSP & PFSP $\times$ $\checkmark$
Dota 2 10 $\times$ OpenAI Five [14] Naive SP & PFSP $\times$ $\times$
Honor of Kings 10 $\times$ MOBA AI [135] $\delta$-uniform SP $\checkmark$ $\checkmark$
Google Research Football 22 $\checkmark$ TiZero [144] FSP & PFSP $\times$ $\times$
Table III: Overview of Representative Studies in Empirical Analysis.
V Open Problems and Future Work
Self-play approaches have demonstrated superior performance due to their unique iterative learning processes and ability to adapt to complex environments. However, there are still several areas that require further research and development.
V-A Theoretical Foundation
Although NE has been shown to exist in games with finite players and finite actions [145], computing NE in larger games remains challenging and consequently, many studies aim to achieve approximate NE [146]. However, in some cases, even computing an approximate NE is difficult [65]. Some research has resorted to higher levels of equilibrium, such as CE [47] and $\alpha$-rank [46].
Although many algorithms are developed with theoretical foundations in game theory, there is often a gap when it comes to applying these algorithms to complex real-world scenarios. For example, although AlphaGo [8], AlphaStar [147] and OpenAI Five [14] have achieved empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by combining empirical success with theoretical validation. This could involve developing new algorithms that are both theoretically sound and practically effective in complex scenarios or proving the theoretical underpinnings of existing successful algorithms in complex environments.
V-B Non-stationarity of the Environment
The strategies of the opponent players evolve as training progresses, and the opponents are a vital factor of the environment in the self-play framework. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling into self-play [121] can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
V-C Scalability and Training Efficiency
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five [14], the hero pool size is limited to only 17 heroes. MOBA AI [135] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is to leverage the inherent connections among players to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments.
These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods, which involve both computational and storage aspects. Computational efficiency is a limiting factor due to the iterative nature of self-play, where agents repeatedly play against themselves or past versions, requiring extensive computational resources. Moreover, although forming more complex populations and competitive mechanisms [147] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. Additionally, self-play is storage-intensive, as it requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which need to store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
V-D With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [148], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning [149, 150] is a key step in aligning LLM with more desired behaviors, but requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [151] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Some other work [152, 153] also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. SPAG [154] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on a wide range of reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero [155] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work [156] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research efforts.
V-E Reality Applications
Self-play is a powerful technique with widespread applications across various domains. It is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks [157]. Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) [158]. Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors [159] and enables vehicles to learn negotiation strategies to merge on or off roads [160], although currently within 2D simulators.
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [161], and to address the image retargeting problem [162]. Moreover, EvoPlay [163] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [164] to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for adversarial catching tasks, with significant efforts dedicated to Sim2Real transitions for real-world success [165].
VI Conclusion
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play, in which agents interact with copies or past versions of themselves, has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Details are meticulously provided on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within complex scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Future work may focus on solving these problems, integrating self-play with LLMs, and achieving real-world applications. In conclusion, self-play stands as a cornerstone of modern RL research, offering profound insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this dynamic and evolving field.
Acknowledgment
This research was supported by National Natural Science Foundation of China (No.62325405, U19B2019, M-0248), Tsinghua University Initiative Scientific Research Program, Tsinghua-Meituan Joint Institute for Digital Life, Beijing National Research Center for Information Science, Technology (BNRist) and Beijing Innovation Center for Future Chips, Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240830, China Postdoctoral Science Special Foundation 2024T170496.